If Veronica Vaughn spends all of her daily income on cigarettes and Yoo Hoo, she can afford 10 packs of cigarettes and 10 bottles of Yoo Hoo. She can also afford 6 packs of cigarettes and 22 bottles of Yoo Hoo.



Draw verconica's daily budget set with packs of cigarettes on the x axis and bottles of yoo-hoo on the y axis

1
Expert's answer
2021-11-12T07:59:55-0500

(a)) I = PxX + PyY

Where: I = Income

Px = Price of cigarettes

Py = Price of Yoo Hoo

X = Cigarettes

Y = Yoo Hoo

I = 10Px + 10Py

I= 6Px + 22Py

I/10 = Px

I/10 = Py

Budget line graph for

I = 10Px + 10Py